Use Laplace transform to solve initial value problem.

Y^2+Y={t,0 less than or equal to t less than 1

{0,t greater or equal to 1

y(0)=0,y'(0)=0


1
Expert's answer
2021-06-22T13:57:29-0400

Given Y'' + Y =  "\begin{cases}
t, \text{ if } 0 \leq t < 1 \\
0, \text{ if } t \geq 1
\end{cases}"


let f(t) = "\begin{cases}
t, \text{ if } 0 \leq t < 1 \\
0, \text{ if } t \geq 1
\end{cases}"

we can express f(t) using unit step function U(t - a) = "\begin{cases}
0, \text{ if } t < a \\
1, \text{ if } t \geq a
\end{cases}"

hence we can express f(t) as f(t) = "t" "-" "t"U(t-1)

now given differential equation becomes, Y'' + Y = "t" "-" "t"U(t-1)

applying Laplace transform both side, ( "\mathcal{L}"(f(t)) = F(s) )

"\mathcal{L}"(Y'' + Y) = "\mathcal{L}""(t" "-" "t"U(t-1))

"\mathcal{L}"(Y'') + "\mathcal{L}" (Y) = "\mathcal{L}"( "t" ) "-" "\mathcal{L}"("t"U(t-1))


s2Y(s) - sY(0) - Y'(0) + Y(s) = "\frac{1}{s^2}" - e-s"\mathcal{L}"(t+1)


s2Y(s) - sY(0) - Y'(0) + Y(s) = "\frac{1}{s^2}" - e-s"\frac{s+1}{s^2}"


since Y(0) = 0 and Y'(0) = 0 (as given)


s2Y(s) + Y(s) = "\frac{1}{s^2}" - e-s"\frac{s+1}{s^2}"


Y(s) = "\frac{1}{s^2 (s^2+1)}" - e-s"\frac{s+1}{s^2(s^2+1)}"


now applying the inverse Laplace transform we get,


"\mathcal{L}^{-1}" Y(s) = "\mathcal{L}^{-1}" ( "\frac{1}{s^2 (s^2+1)}" - e-s"\frac{s+1}{s^2(s^2+1)}" )


Y(t) = "\mathcal{L}^{-1}" ( "\frac{1}{s^2 (s^2+1)}") -"\mathcal{L}^{-1}" (e-s"\frac{s+1}{s^2(s^2+1)}" )


Y(t) = "\mathcal{L}^{-1}" ( "\frac{1}{s^2 } - \frac{1}{(s^2+1)}") -"\mathcal{L}^{-1}" (e-s"\frac{s+1}{s^2(s^2+1)}" )


Y(t) = "\mathcal{L}^{-1}" ( "\frac{1}{s^2 } - \frac{1}{(s^2+1)}") -"\mathcal{L}^{-1}" (e-s"\frac{s+1}{s^2(s^2+1)}" )


Y(t) = "\mathcal{L}^{-1}" "(\frac{1}{s^2 }) - \mathcal{L}^{-1}(\frac{1}{(s^2+1)})" "-\mathcal{L}^{-1}" (e-s"\frac{1}{s^2 (s^2+1)}"+ e-s "\frac{s}{s^2 (s^2+1)}")


Y(t) = "\mathcal{L}^{-1}" "(\frac{1}{s^2 }) - \mathcal{L}^{-1}(\frac{1}{(s^2+1)})""+\mathcal{L}^{-1}" e-s"(\frac{1}{s^2 })" "-" "\mathcal{L}^{-1}" e-s "(\frac{1}{(s^2+1)})" +"\mathcal{L}^{-1}" e-s"(\frac{1}{s })" "-" "\mathcal{L}^{-1}" e-s "(\frac{s}{(s^2+1)})"


Y(t) = t - sin(t) + tU(t-1) -sin(t-1)U(t-1) - cos(t-1)U(t-1) +U(t-1)


Y(t) = t - sin(t) + (t -sin(t-1) - cos(t-1) +1)U(t-1) where U(t-1) is the unit step function.